OpenStudy (anonymous):

The scatter plot shows the number of pages scanned (y) in different number of hours (x) by a scanning machine:

OpenStudy (anonymous):

Which function best represents the data shown in the scatter plot?

OpenStudy (anonymous):

y = 100 + 1,000x y = 100 - 1,000x y = 500 - 1,000x y = 500 + 1,000x

OpenStudy (anonymous):

y=500+1,000x by simply put x=0,1,2,3